how many buckets of capacity 15 litre can be filled from a water tank 4m long, 2m wide and 1.2m high
Answer by srinivas.g(540)

capacity of te water tank = length x width x hight
= 4 x2 x 1.2
= 96 m^3
but 1 m^3 =1000 liters
= 96 *1000
= 96000 liters
Capacity of tank = 96000 liters
capacity of each bucket = 15 liters
no of buckets needed to fill = total capacity of water tank / capacity of each bucket
= 96000/15
=6400
Result :6400